If a/b=b/10=c=4 then a+b+c/c=?
Answer by Alan3354(69443)

If a/b=b/10=c=4 then a+b+c/c=?
-----------------
c=4
--
b/10 = 4
b = 40
--
a/b = 4
a/40 = 4
a = 160
==================
(160 + 40 + 4)/4 = 51